# Relative monodromy of abelian logarithms for finite covers of universal families

### v2

@@ Abstract
Let’s fix a complex abelian scheme endowed with a non-torsion section and with a finite surjective modular map onto a universal family of abelian varieties (with some fixed $-\ell$-structure and without locally constant parts). We show that the relative monodromy group of the abelian logarithm is non-trivial and of full rank. As a consequence we deduce a new proof of Manin’s kernel theorem and of the algebraic independence of the coordinates of abelian logarithms with respect to the coordinates of periods.
@@ 1 Introduction
A $g$-dimensional abelian scheme $\phi\colon\mathcal{A}\to S$ over $\mathbb{C}$ can be seen as family of $g$-dimensional complex abelian varieties $\{\mathcal{A}_{s}\}_{s\in S(\mathbb{C})}$ (the fibres of $\phi$) that “varies algebraically accordingly to $\phi$” over the parameterizing space $S(\mathbb{C})$. Each fibre $\mathcal{A}_{s}$ can be analytically identified with a complex torus $\mathbb{C}^{g}/\Lambda_{s}$. Locally, on simply connected subsets $U\subset S(\mathbb{C})$ it is possible to define a holomorphic period map $\mathfrak{P}:U\to\operatorname{Lie}(\mathcal{A})^{2g}$ where $\operatorname{Lie}(\mathcal{A})$ is the complex lie algebra bundle of $\mathcal{A}(\mathbb{C})$. The period map associates to any point $s\in U$ a $\mathbb{R}$-basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$ and in general it cannot be globally defined on $S(\mathbb{C})$. The obstruction to the existence of the analytic prolongation of $\mathfrak{P}$ along a loop around $s\in S(\mathbb{C})$ is measured by the monodromy group $\textrm{Mon}(\mathcal{A},s)$. Given an algebraic section of the abelian scheme, which is a $\mathbb{C}$-morphism $\sigma\colon S\to\mathcal{A}$, again locally on simply connected open subsets $U^{\prime}\subset S(\mathbb{C})$ it is possible to lift $\sigma$ through the exponential map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$ to a holomorphic map $\log_{\sigma}\colon U^{\prime}\to\operatorname{Lie}(\mathcal{A})$. This map is called an abelian logarithm of the section $\sigma$. On a common open set $V\subset S(\mathbb{C})$ where both $\log_{\sigma}$ and $\mathfrak{P}$ are defined, we can express $\log_{\sigma}(s)\in\mathbb{C}^{g}$ in terms of the basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$. The coordinates form a real-analytic function $\beta_{\sigma}\colon V\to\mathbb{R}^{2g}$ which is nowadays called the Betti map (associated to $\sigma$) and that turned out to be of crucial importance in the theory of unlikely intersections. Such map was implicitly considered by Manin in [23], to get a proof of Mordell’s conjecture in the function field case. The aforementioned abelian logarithm $\log_{\sigma}$ is in general defined only locally, thus it is interesting to measure the obstruction to the existence of a global abelian logarithm: the simultaneous monodromy of the period map and abelian logarithm of $\sigma$ gives rise to the monodromy group of the section $M_{\sigma}$. The knowledge of $\textrm{Mon}(\mathcal{A},s)$ and the monodromy group of the logarithm of a section gives relevant information on the structure of the abelian scheme and on the section, respectively: in fact, no non-zero periods can be globally defined when the abelian scheme has no fixed part (see [15, Lemma 5.6]) and the monodromy of a logarithm is trivial only when the section is the zero-section of the scheme. It is usually important to understand more on the structure of monodromy groups. In [8] an [26] the authors construct examples where the monodromy group $\textrm{Mon}(\mathcal{A},s)$ is of infinite index in the arithmetic group attached to its Zariski closure, i.e. thin. Also in [30] it is pointed out that thin groups often arise in the context of monodromy groups. Given a finitely generated group in $\textrm{GL}_{n}(\mathbb{Z})$, usually it’s not too difficult to compute its Zariski closure, but deciding if the group is thin can be very subtle and difficult. Thus, in general it is very interesting to determine under which hypotheses the monodromy groups are big in their Zariski closures.
@@
We investigate the obstruction to have a globally defined abelian logarithm along loops inducing a trivial monodromy action on the period map. This obstruction is measured by the relative monodromy group $M^{\textrm{rel}}_{\sigma}$. It is easy to check that $M^{\textrm{rel}}_{\sigma}$ is a $\mathbb{Z}$-submodule of $\mathbb{Z}^{2g}$, so it is interesting to understand how its rank is related to $\sigma$. The group $M_{\sigma}^{\textrm{rel}}$ gives relevant information on the section, in particular its non-triviality should be related to the property of $\sigma$ of being non-torsion, and its rank should give information on the minimal group subscheme containing the image of $\sigma$.
@@
In the case of non-isotrivial elliptic schemes Corvaja and Zannier proved in [9] that when $\sigma$ is a non-torsion section $M^{\textrm{rel}}_{\sigma}$ is nontrivial and moreover it is isomorphic to $\mathbb{Z}^{2}$ (so it has full rank). In the same paper, they provided an example showing that the relative monodromy group could be smaller than expected (even trivial) when compared to the Zariski closures of non-relative monodromy groups: therefore, there are some groups which cannot arise as monodromy groups in this setting. Moreover, we point out that the second author of the present paper in [33] studied the relative monodromy for products of two elliptic schemes. In [34] he also gives an explicit proof of the results of Corvaja and Zannier.
In general it is conjectured that for any abelian scheme $\phi\colon\mathcal{A}\to S$ without fixed part the rank of $M^{\textrm{rel}}_{\sigma}$ is $2g^{\prime}$ where $g^{\prime}$ is the relative dimension of the minimal group subscheme of $\mathcal{A}\to S$ containing the image of the section $\sigma$. In this paper we give a positive answer to this conjecture in the case of an abelian scheme admitting a finite base change endowed with a surjective finite map onto a universal family $\mathfrak{A}_{g}\to\mathbb{A}_{g}$ with some fixed $-\ell$-structure and without locally constant parts. Specifically we prove the following bigness results about relative monodromy groups:
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme such that (up to a finite base change) the modular map $p:S\to\mathbb{A}_{g}$ is a finite surjective morphism. If $\sigma:S\to\mathcal{A}$ is a non-torsion section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is non-trivial.
Theorem 1.2.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme such that (up to a finite base change) the modular map $p:S\to\mathbb{A}_{g}$ is a finite surjective morphism. If $\sigma:S\to\mathcal{A}$ is a non-torsion section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
Notice that in our statements we don’t need to distinguish the cases when the image of the section $\sigma$ is contained in proper group subschemes since having a finite map onto a universal family prevents our abelian scheme from having any non-trivial proper subscheme.
@@
These theorems have immediate implications on the arithmetic properties of the sections of abelian schemes, in particular on some properties of the Betti map. On one hand periods and abelian logarithms in general cannot be globally defined, but on the other hand it’s also true that they can be always be globally defined after going to the universal cover $\widetilde{S}$ of the parameter base $S$, so the Betti map is also globally defined on $\widetilde{S}$. Let us denote by $S_{*}\to S$ the minimal unramified cover where periods are globally defined and by $S_{\sigma}\to S_{*}$ the minimal unramified cover where abelian logarithm of $\sigma$ is globally defined: then, $S_{\sigma}\to S_{*}$ is also the minimal unramified cover of $S_{*}$ where the Betti map is globally defined and we have the following situation
@@
The map $S_{*}\to S$ is a Galois cover whose Galois group is $\textrm{Mon}(\mathcal{A},s)$, while the map $S_{\sigma}\to S$ is a Galois cover with Galois group $M_{\sigma}$. Our main theorems determine properties of the Galois group $M_{\sigma}^{\textrm{rel}}$ of the Galois cover $S_{\sigma}\to S_{*}$, giving precise information about the global definition of the Betti map which is useful to know for several applications.
For instance, it is straightforward to see that if $\sigma$ is a torsion section then the Betti map of $\sigma$ is constant and thus globally defined on the base. The converse statement is also true but more difficult to prove, it is widely known as Manin’s kernel theorem. As a consequence of 1.1, we obtain a new proof of a stronger version of Manin’s kernel theorem (under our hypotheses). We call it “strong version” because in its statement we just assume the Betti coordinates to be globally defined instead of rational constant a priori.
@@
The Betti map is used in the work [28] of Pila and Zannier together with a result of Pila and Wilkie in transcendental Diophantine geometry (see [27]) to give a new proof of the Manin-Mumford conjecture. Such a strategy is referred as “the Pila-Zannier method”; in the case of abelian schemes over $\overline{\mathbb{Q}}$ it works in the following way: the goal is to have a control on the distribution of torsion values of $\sigma$ (i.e. elements of $\sigma^{-1}(\mathcal{A}_{\operatorname{tor}})$) with bounded heights. First one gives lower bounds on the number of such torsion values using Galois conjugates and the height inequality of Dimitrov-Gao-Habegger. On the other hand the Betti map $\beta_{\sigma}$ transforms the torsion values in rational points of a definable set in the $o$-minimal structure $\mathbb{R}_{\operatorname{an},\exp}$. At this point a “Pila-Wilkie type” result gives a control on the rational points on the transcendental part of such definable set. Still we have no information on the rational points in the algebraic part of the definable set, but here the crucial point is to use some functional transcendence theorem of “Ax-Schanuel type” to control the algebraic part of the definable set. Let’s take a closer look at the last step involving the functional transcendence arguments: a classical way to tackle this step is by using a result of André (see [1, Theorem 3]). It says that the coordinates of the abelian logarithm $\log_{\sigma}$ are pairwise algebraically independent over the extension of $\mathbb{C}(S)$ generated by the coordinates of the period map $\mathfrak{P}$. Such result has often very strong implications, for instance in [13], thanks to such result, it is shown that the algebraic part of the definable set constructed with the Pila-Zannier method is empty.
@@
is equal to the dimension of the kernel of the homomorphism from the differential Galois group of $\mathbb{C}(S)(\mathfrak{P},\partial\mathfrak{P})(\log_{\sigma},\partial\log_{\sigma})/\mathbb{C}(S)$ to the differential Galois group of $\mathbb{C}(S)(\mathfrak{P},\partial\mathfrak{P})/\mathbb{C}(S)$. Recall that by general theory, the differential Galois group of a Picard-Vessiot extension is exactly the Zariski closure of the corresponding monodromy group. In [9] it is pointed out that André’s theorem gives no information about the relative monodromy group $M_{\sigma}^{\textrm{rel}}$. Thus, 1.2 can be interpreted as a strengthening of André’s theorem (under our hypotheses). We shall in fact prove how André’s result is a consequence of 1.2, and this in turns leads to a new proof of the algebraic independence of the coordinates of $\log_{\sigma}$ with respect to periods.
@@
Our hope for future developments is twofold. First of all, the main theorems of this paper could be applied for the solution of some “special points problems” when in the functional transcendence step of the Pila-Zannier method André’s theorem fails to give a control on rational points of the algebraic part of the constructed definable set. It is is not clear yet if such cases exist and if they are interesting. Secondly, it would be nice to remove the main hypothesis on the surjectivity of the modular map $p\colon S\to\mathbb{A}_{g}$. Such assumption is necessary in order to apply the generalization of Shioda’s theorem due to Mok and To (see [25, Main Theorem]). This result has in fact been proved only for Kuga families. One could think about a different proof of our main theorems that doesn’t invoke [25], but at the moment the required strategy is not entirely understood.
@@
The approach of our proofs was originally inspired by [9], and it is based on the interpretation of the obstruction to globally define a logarithm as a Galois cohomology class with respect to the monodromy action of the fundamental group of $S(\mathbb{C})$ on $\mathbb{Z}^{2g}$. The new key ideas consist of using the Lefschetz hyperplane theorem for quasi-projective varieties (see [18, 2.2]) and a theorem of Mok and To (see [25, Main Theorem]). In the first step of the proof of 1.1 we construct an intermediate abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ which factorizes the universal diagram induced by the modular map $p:S\to\mathbb{A}_{g}$. This construction is carried out by means of topological techniques in such a way that $\mathcal{A}^{\prime}\to S^{\prime}$ is a finite unramified base change of the universal family (a Kuga family). In the second step we find a suitable curve inside $S(\mathbb{C})$ which preserves the complete monodromy action of the fundamental group of $S(\mathbb{C})$ on periods and logarithms. The curve is obtained after iterating the aforementioned Lefschetz theorem and turns out to be a Stein space. This last property allows us to construct holomorphic sections of the Kuga family coming from a Galois cohomology class describing the obstruction of the global logarithms. Then the proof follows by a contradiction argument: we assume the triviality of the relative monodromy group for a non-torsion section $\sigma$, then by using the previous apparatus and the properties of trace and pull-back operators, we deduce that $\sigma$ has a global logarithm. This finally contradicts Lang-Néron theorem on the Mordell-Weil group of the generic fiber. This proves the non-triviality of $M_{\sigma}^{\textrm{rel}}$.
In order to prove 1.2 on the maximality of the rank of $M_{\sigma}^{\textrm{rel}}$ the main key ingredient is to exploit the irreducibility of the monodromy action on the lattice of periods.
@@
In Section 2 we define all the main objects and we collect the results coming from the general theory. In particular, we give a detailed account of the various monodromy groups. In Section 3 we carry out the main proofs after some technical results about “the invariance” of our main theorems. Section 4 is devoted to some useful applications of our results.
Acknowledgments.
The authors are deeply grateful to Y. André, P. Corvaja and U. Zannier for the useful discussions on the topic. The second author is supported by PRIN 2022, CUP F53D23002740006. Moreover, he would like to thank the Institute for Theoretical Sciences of Westlake University for their hospitality while he was completing the paper.

### v3

# Relative monodromy of ramified sections on abelian schemes

@@ Abstract
Let’s fix a complex abelian scheme $\mathcal{A}\to S$ of relative dimension $g$, without fixed part, and having maximal variation in moduli. We show that the relative monodromy group $M^{\textrm{rel}}_{\sigma}$ of a ramified section $\sigma\colon S\to\mathcal{A}$ is nontrivial. Moreover, under some hypotheses on the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ we show that $M^{\textrm{rel}}_{\sigma}\cong\mathbb{Z}^{2g}$. We discuss several examples and applications. For instance we provide a new proof of Manin’s kernel theorem and of the algebraic independence of the coordinates of abelian logarithms with respect to the coordinates of periods.
@@ 1 Introduction
Let $S$ be a regular, irreducible, quasi-projective variety defined over $\mathbb{C}$. An abelian scheme $\phi\colon\mathcal{A}\to S$ over $\mathbb{C}$ of relative dimension $g$ can be seen as family of $g$-dimensional complex abelian varieties $\{\mathcal{A}_{s}\}_{s\in S(\mathbb{C})}$ (the fibres of $\phi$) that “varies algebraically accordingly to $\phi$” over the parameterizing space $S(\mathbb{C})$. Each fibre $\mathcal{A}_{s}$ can be analytically identified with a complex torus $\mathbb{C}^{g}/\Lambda_{s}$. Locally, on simply connected subsets $U\subset S(\mathbb{C})$ it is possible to define a holomorphic period map $\mathfrak{P}:U\to\operatorname{Lie}(\mathcal{A})^{2g}$ where $\operatorname{Lie}(\mathcal{A})$ is the complex Lie algebra bundle of $\mathcal{A}(\mathbb{C})$. The period map associates to any point $s\in U$ a $\mathbb{R}$-basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$ and in general it cannot be globally defined on $S(\mathbb{C})$. The obstruction on the existence of the analytic continuation of $\mathfrak{P}$ along a loop around $s\in S(\mathbb{C})$ is measured by the monodromy group $\textrm{Mon}(\mathcal{A},s)$. Given an algebraic section of the abelian scheme, which is a $\mathbb{C}$-morphism $\sigma\colon S\to\mathcal{A}$, again locally on simply connected open subsets $U^{\prime}\subset S(\mathbb{C})$ it is possible to lift $\sigma$ through the exponential map $\exp\colon\operatorname{Lie}(\mathcal{A})\to\mathcal{A}(\mathbb{C})$ to a holomorphic map $\log_{\sigma}\colon U^{\prime}\to\operatorname{Lie}(\mathcal{A})$. This map is called an abelian logarithm of the section $\sigma$. On a common open set $V\subset S(\mathbb{C})$ where both $\log_{\sigma}$ and $\mathfrak{P}$ are defined, we can express $\log_{\sigma}(s)\in\mathbb{C}^{g}$ in terms of the basis $\mathfrak{P}(s)$ of the lattice $\Lambda_{s}$. The coordinates form a real-analytic function $\beta_{\sigma}\colon V\to\mathbb{R}^{2g}$ which is nowadays called the Betti map (associated to $\sigma$) and that turned out to be of crucial importance in the theory of unlikely intersections. Such map was implicitly considered by Manin in [25], to get a proof of Mordell’s conjecture in the function field case. The aforementioned abelian logarithm $\log_{\sigma}$ is in general defined only locally, thus it is interesting to measure the obstruction to the existence of a global abelian logarithm: the simultaneous monodromy of the period map and abelian logarithm of $\sigma$ gives rise to the monodromy group of the section $M_{\sigma}$. The knowledge of $\textrm{Mon}(\mathcal{A},s)$ and the monodromy group of the logarithm of a section gives relevant information on the structure of the abelian scheme and on the section, respectively: in fact, no non-zero periods can be globally defined when the abelian scheme has no fixed part (see [16, Lemma 5.6]) and the monodromy of a logarithm is trivial only when the section is the zero-section of the scheme. It is usually important to understand more on the structure of monodromy groups. In [9] an [27] the authors construct examples where the monodromy group $\textrm{Mon}(\mathcal{A},s)$ is of infinite index in the arithmetic group attached to its Zariski closure, i.e. thin. Also in [30] it is pointed out that thin groups often arise in the context of monodromy groups. Given a finitely generated group in $\textrm{GL}_{n}(\mathbb{Z})$, usually it’s not too difficult to compute its Zariski closure, but deciding if the group is thin can be very subtle and difficult. Thus, in general it is very interesting to determine under which hypotheses the monodromy groups are big in their Zariski closures.
@@
We investigate the obstruction that prevents the existence of a globally defined abelian logarithm along loops which induce a trivial monodromy action on the period map. This obstruction is measured by the relative monodromy group $M^{\textrm{rel}}_{\sigma}$. It is easy to check that $M^{\textrm{rel}}_{\sigma}$ is a $\mathbb{Z}$-submodule of $\mathbb{Z}^{2g}$, so it is interesting to understand how its rank is related to $\sigma$. The group $M_{\sigma}^{\textrm{rel}}$ gives relevant information on the section, in particular its non-triviality should be related to the property of $\sigma$ of being non-torsion, and its rank should give information on the minimal group subscheme containing the image of $\sigma$.
@@
In the case of non-isotrivial elliptic schemes Corvaja and Zannier proved in [10] that when $\sigma$ is a non-torsion section the group $M^{\textrm{rel}}_{\sigma}$ is non-trivial and moreover it is isomorphic to $\mathbb{Z}^{2}$ (so it has full rank). In the same paper, they provided an example showing that a priori, if we just limit ourselves to group theoretic arguments, the relative monodromy group could be smaller than expected (even trivial) when compared to the Zariski closures of non-relative monodromy groups. Hence, in this setting there are some groups which cannot arise as monodromy groups of sections. Moreover, we point out that the second author of the present paper in [33] studied the relative monodromy for products of two elliptic schemes. In [32] he also gives an explicit proof of the results of Corvaja and Zannier.
In this paper we generalize the work of [10] on the relative monodromy $M^{\textrm{rel}}_{\sigma}$, focusing on ramified sections instead of non-torsion sections. The notion of ramified section is formally introduced below but it essentially means that $\sigma:S\to\mathcal{A}$ doesn’t come from a base-change of an abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ whose modular map is unramified.
Let $\mathfrak{A}_{g,\ell}\to\mathbb{A}_{g,\ell}$ be a universal family with some fixed $-\ell$-structure and without locally constant parts; in order to simplify the notation, in the rest of the paper we drop the subscript $\ell$ and simply denote the family by $\mathfrak{A}_{g}\to\mathbb{A}_{g}$. Moreover, in order to make the following definition meaningful, we also need to assume that the base $S$ is maximal in the sense that on any compactification $\overline{A}\to\overline{S}$ the fibers over $\overline{S}\setminus S$ are all singular.
Definition 1.1.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme with modular map $p:S\to\mathbb{A}_{g}$. We say that a section $\sigma:S\to\mathcal{A}$ is unramified (over $\mathbb{A}_{g}$) if:
  • $(i)$
    There exist an unramified morphism $q^{\prime}:S^{\prime}\to p(S)$ and an abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ such that $\mathcal{A}=\mathcal{A}^{\prime}\times_{S^{\prime}}S$ and the following diagram is commutative:
[…]
  • $(ii)$
    There exists a section $\sigma^{\prime}:S^{\prime}\to\mathcal{A}^{\prime}$ with the following property: let $q^{*}\sigma^{\prime}\colon S\to\mathcal{A}$ be the natural section induced by $\sigma^{\prime}\circ q$, then $\sigma=q^{*}\sigma^{\prime}$.
A section $\sigma\colon S\to\mathcal{A}$ is said to be ramified (over $\mathbb{A}_{g}$) if it is not unramified.
Under the extra hypotheses that the abelian scheme is without fixed part (i.e. it has trivial Chow trace) and that the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite [^1] In the literature the is no uniform consensus on the precise notion of generically finite morphism. For us a morphism of schemes $f\colon X\to Y$ is generically finite if there exists a dense open $V\subseteq Y$ such that the restriction $f\colon f^{-1}(V)\to V$ is a finite morphism. (onto its image) we prove the result stated below. Notice that such condition on the modular map is often referred as maximal variation in moduli.
@@
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is non-trivial.
Our techniques seem unsuitable to prove the same statement for unramified non-torsion sections. However, when it is possible to apply [26, Main Theorem] on the finiteness of the group of holomorphic sections for Kuga families, we extend 1.2 for all non-torsion sections. In particular, this happens when the modular map is generically finite onto a weakly-special subvariety of $\mathbb{A}_{g}$.
In addition, we provide a precise result on the size of $M^{\textrm{rel}}_{\sigma}$ in some special cases:
Corollary 1.3.
Let $\phi:\mathcal{A}\to S$ be an abelian scheme without fixed part such that (up to a finite base change) the modular map $p:S\to p(S)\subseteq\mathbb{A}_{g}$ is generically finite. If $\sigma:S\to\mathcal{A}$ is a ramified section and the action of the monodromy group $\textrm{Mon}(\mathcal{A})$ is irreducible, then the relative monodromy group $M_{\sigma}^{\textrm{rel}}$ is isomorphic to $\mathbb{Z}^{2g}$.
In a similar way as above 1.3 can be extended for all non-torsion sections under some additional hypothesis. Note that the action of $\textrm{Mon}(\mathcal{A})$ is irreducible for instance when $p(S)$ is not contained in any proper special subvariety of $\mathbb{A}_{g}$.
1.3 partially settles the following general conjecture:
Conjecture 1.4.
Let $\phi\colon\mathcal{A}\to S$ be an abelian scheme of relative dimension $g$ without fixed part. The rank of $M^{\textrm{rel}}_{\sigma}$ is $2g^{\prime}$ where $g^{\prime}\leq g$ is the relative dimension of the minimal group subscheme of $\mathcal{A}\to S$ containing the image of the section $\sigma$.
After the first version of this paper (which contained slightly weaker results) was made public, we privately received the preprint [2] of Y. André who partially addresses 1.4, including some of our cases. Under certain hypotheses, including the requirement that the monodromony of periods is not thin, he shows that $M^{\textrm{rel}}_{\sigma}$ is of maximal rank. See Section 3.5 for more details and an explicit example in which the monodromy of periods is thin and $M^{\textrm{rel}}_{\sigma}$ is of maximal rank.
@@
Our theorem has immediate implications on the arithmetic properties of the sections of abelian schemes, in particular on some properties of the Betti map. On one hand periods and abelian logarithms in general cannot be globally defined, but on the other hand it is also true that they can be always be globally defined after going to the universal cover $\widetilde{S}$ of the parameter base $S$, so the Betti map is also globally defined on $\widetilde{S}$. Let us denote by $S_{*}\to S$ the minimal unramified cover where periods are globally defined and by $S_{\sigma}\to S_{*}$ the minimal unramified cover where abelian logarithm of $\sigma$ is globally defined: then, $S_{\sigma}\to S_{*}$ is also the minimal unramified cover of $S_{*}$ where the Betti map is globally defined and we have the following situation
@@
The map $S_{*}\to S$ is a Galois cover whose Galois group is $\textrm{Mon}(\mathcal{A},s)$, while the map $S_{\sigma}\to S$ is a Galois cover with Galois group $M_{\sigma}$. Our main theorem determines properties of the Galois group $M_{\sigma}^{\textrm{rel}}$ of the Galois cover $S_{\sigma}\to S_{*}$, giving precise information about the global definition of the Betti map which is useful to know for several applications.
For instance, it is straightforward to see that if $\sigma$ is a torsion section then the Betti map of $\sigma$ is constant and thus globally defined on the base. The converse statement is also true but more difficult to prove, it is widely known as Manin’s kernel theorem. As a consequence of 1.2, we obtain a new proof of a stronger version of Manin’s kernel theorem. We call it “strong version” because in its statement we just assume the Betti coordinates to be globally defined instead of rational constant a priori.
@@
The Betti map is used in the work [29] of Pila and Zannier together with a result of Pila and Wilkie in transcendental Diophantine geometry (see [28]) to give a new proof of the Manin-Mumford conjecture. Such a strategy is referred as “the Pila-Zannier method”; in the case of abelian schemes over $\overline{\mathbb{Q}}$ it works in the following way: the goal is to have a control on the distribution of torsion values of $\sigma$ (i.e. elements of $\sigma^{-1}(\mathcal{A}_{\operatorname{tor}})$) with bounded heights. First one gives lower bounds on the number of such torsion values using Galois conjugates and the height inequality of Dimitrov-Gao-Habegger. On the other hand the Betti map $\beta_{\sigma}$ transforms the torsion values in rational points of a definable set in the $o$-minimal structure $\mathbb{R}_{\operatorname{an},\exp}$. At this point a “Pila-Wilkie type” result gives a control on the rational points on the transcendental part of such definable set. Still we have no information on the rational points in the algebraic part of the definable set, but here the crucial point is to use some functional transcendence theorem of “Ax-Schanuel type” to control the algebraic part of the definable set. Let us take a closer look at the last step involving the functional transcendence arguments: a classical way to tackle this step is by using a result of André (see [1, Theorem 3]). It says that the coordinates of the abelian logarithm $\log_{\sigma}$ are pairwise algebraically independent over the extension of $\mathbb{C}(S)$ generated by the coordinates of the period map $\mathfrak{P}$. Such result has often very strong implications, for instance in [14], thanks to such result, it is shown that the algebraic part of the definable set constructed with the Pila-Zannier method is empty.
@@
is equal to the dimension of the kernel of the homomorphism from the differential Galois group of $\mathbb{C}(S)(\mathfrak{P},\partial\mathfrak{P})(\log_{\sigma},\partial\log_{\sigma})/\mathbb{C}(S)$ to the differential Galois group of $\mathbb{C}(S)(\mathfrak{P},\partial\mathfrak{P})/\mathbb{C}(S)$. Recall that by general theory, the differential Galois group of a Picard-Vessiot extension is exactly the Zariski closure of the corresponding monodromy group. In [10] it is pointed out that André’s theorem gives no information about the relative monodromy group $M_{\sigma}^{\textrm{rel}}$. Thus, 1.3 can be interpreted as a strengthening of André’s theorem (under our hypothesis). We shall in fact prove how André’s result is a consequence of 1.3, and this in turns leads to a new proof of the algebraic independence of the coordinates of $\log_{\sigma}$ with respect to periods.
@@
Our hope for future developments is twofold. First of all, the main results of this paper could be applied for the solution of some “special points problems” when in the functional transcendence step of the Pila-Zannier method André’s theorem fails to give a control on rational points of the algebraic part of the constructed definable set. It is not clear yet if such cases exist and if they are interesting. Secondly, it would be nice to remove the hypothesis of maximal variation in moduli and also to obtain the result for all non-torsion sections in 1.2.
@@
The approach of our proofs was originally inspired by [10], and it is based on the interpretation of the obstruction to globally define a logarithm as a Galois cohomology class with respect to the monodromy action of the fundamental group of $S(\mathbb{C})$ on $\mathbb{Z}^{2g}$. The new key ideas consist of using the Lefschetz hyperplane theorem for quasi-projective varieties (see [19, 2.2]) to restrict our reasoning to generic curves on the base. At this point we exploit the properties of the trace operator with respect to the pullback of the section in the case of finite morphisms, but as opposed to [10] we don’t need to construct abelian schemes with finite Mordell-Weil group. [^2] In a previous version of the paper we considered only Kuga families so that we could apply [26, Main Theorem].
For the proof of 1.2 we proceed in the following way: we construct an intermediate abelian scheme $\mathcal{A}^{\prime}\to S^{\prime}$ which factorizes the universal diagram induced by the modular map $p:S\to\mathbb{A}_{g}$. This construction is carried out by means of topological techniques in such a way that $\mathcal{A}^{\prime}\to S^{\prime}$ is a finite unramified base change of the universal family. Then we find a suitable curve inside $S(\mathbb{C})$ which preserves the complete monodromy action of the fundamental group of $S(\mathbb{C})$ on periods and logarithms. The curve is obtained after iterating the aforementioned Lefschetz theorem and turns out to be a Stein space. This last property allows us to construct holomorphic sections of the unramified family which come from a Galois cohomology class describing the obstruction of the global logarithms. The proof then follows by a contradiction argument: we assume the triviality of $M_{\sigma}^{\textrm{rel}}$, then by using the previous apparatus and the properties of trace and pull-back operators, we deduce that $\sigma$ should be unramified. This proves the non-triviality of $M_{\sigma}^{\textrm{rel}}$.
Once that we have addressed the non-triviality of $M^{\textrm{rel}}_{\sigma}$, the corollaries follow immediately from the extra assumptions.
@@
In Section 2 we define all the main objects and we collect the results coming from the general theory. In particular, we give a detailed account of the various monodromy groups. In Section 3.2 we carry out the proof of 1.2 after some technical results about “the invariance” of our main theorem (Section 3.1). Section 3.3 is devoted to the proof of the corollary. In Section 3.4 and Section 3.5 we provide several examples of abelian schemes with sections for which we are able to estimate the relative monodromy. Finally, Section 4 is devoted to some useful applications of our results.
Acknowledgments and remarks.
The authors are deeply grateful to P. Corvaja and U. Zannier for their support and numerous insights. Moreover, they thank Y. André for the interesting discussions on the topic. The second author is supported by PRIN 2022, CUP F53D23002740006. Moreover, he would like to thank the Institute for Theoretical Sciences of Westlake University for their hospitality while he was completing the paper.